OpenStudy (anonymous):

Integrate (((2y+1)/4) + y )^8 ? see comments for this written better.

OpenStudy (anonymous):

\[\left( \frac{ (2y+1) }{ 4 }+y \right)^{9}\]

OpenStudy (anonymous):

with respect to y

OpenStudy (anonymous):

change it to: \[\int\limits_{}^{}\left( \frac{ 6y + 1 }{ 4 } \right)^n dy\]

hartnn (hartnn):

another way is to substitute the quantity inside the bracket as 'u' then find du=...

OpenStudy (anonymous):

ya, do that. i think i made a mistake ^_^

OpenStudy (anonymous):

let ((6y + 1)/4) = u

OpenStudy (anonymous):

I hate to sound dumb but how did you get from ((2y+1)/4) +y to (6y+1)/4?

OpenStudy (anonymous):

add the fractions. they had to have same denominator so: \[\frac{ 2y + 1 }{ 4 } + y = \frac{ 2y+1 }{ 4 } + y*\frac{ 4 }{4 } = \frac{ 2y + 1 }{ 4 } + \frac{ 4y }{ 4 } = \frac{ 6y+1 }{ 4 }\]

hartnn (hartnn):

why do i feel that is not your initial question ? if so whats you main question ?

OpenStudy (anonymous):

of course (slapping forehead) my initial question i had a typo.

OpenStudy (anonymous):

hehe :P

OpenStudy (anonymous):